Suppose that 3% of computer chips produced by a certain machine are defective. The chips are put into packages of 20 chips for distribution to retailers. What is the probability that a randomly selected package of chips will contain at least 2 defective chips?
1
Expert's answer
2020-05-14T16:50:38-0400

This is a binomial distribution

P(defective) =p=0.03

q=1-p=1-0.03

n=20

Let x represent the number of defective chips.

Probability that at least 2 are defective

"P(x\geq 2)"

P(X=x) ="{n\choose x} p^x q^{n-x}"

"P(X \geq 2)=\sum_{x=2}^{20} 0.03^x 0.97^{20-x}"

=0.1198